Madisonp:

Steve opens a bank account with a simple annual interest rate of 5%. After four year, how much interest will Steve earn on an initial deposit of $800? $160 $140 $60 $40

xLivvx:

The formula of the simple interest is I = Prt, where P is the initial deposit r is the annual rate in decimal t is the time of investment

xLivvx:

Steve opens a bank account with a simple annual interest rate of 5% r = 5% = 5 ÷ 100 = ?

xLivvx:

@Madisonp What would r = 5% = 5 ÷ 100 = ?

xLivvx:

Are u still there?

Madisonp:

yes

xLivvx:

Ok what would r = 5% = 5 ÷ 100 equal ?

Madisonp:

i don't know

Madisonp:

160

xLivvx:

160 is the answer to the question you posted

Madisonp:

is it right

xLivvx:

Yes

Madisonp:

Thank You

Madisonp:

I gave you best response

Madisonp:

YOU WELCOME